OpenStudy (anonymous):

1) find d^2f/dx^2 when f(x) = x e^-7y+5y 2) find d^2f/dx dy when f(x) = x e^-8y+3y

OpenStudy (anonymous):

two different questions here question (1) find d^2f/dx^2when f(x) = x e^-7y+5y question (2) find d^2f/dx dy when f(x) = x e^-8y+3y

OpenStudy (amistre64):

deriveing twice I assume

OpenStudy (amistre64):

do we do partials? or implicits?

OpenStudy (amistre64):

well, since f(x) implies one variable id conjecture implicits

OpenStudy (dumbcow):

hmm since its wrt x and both are linear wrt x then the 2nd derivatives are 0

OpenStudy (amistre64):

\[ x e^{-7y} + 5y\] by chance?

OpenStudy (dumbcow):

oh i was assuming partial

OpenStudy (amistre64):

those groupings are a pain if you dont get them right

OpenStudy (amistre64):

x e^(-7y) + 5y x' e^(-7y)+ x e^(-7y)' + 5 y' e^(-7y) -7x e^(-7y) y' + 5 y' then again

OpenStudy (anonymous):

i have loads of these, need to be handed in ten minutes haha

OpenStudy (amistre64):

the second one there suggests partials but its only a single variable

OpenStudy (dumbcow):

amistre, you are finding dy/dx not df/dx though

OpenStudy (amistre64):

f(x) assumes that all variables are functions of x, so y = y(x)

OpenStudy (amistre64):

the 2nd on tho says f(x) and wants f_xy

OpenStudy (dumbcow):

i think it should be written f(x,y) otherwise it would have asked for dy^2/dx right?

OpenStudy (amistre64):

i think there might be some confusion as to whether f(x) IS y rather than written out is a rule with x(x) and y(x)

OpenStudy (amistre64):

hard for me to parse the intricate details tho :)

OpenStudy (dumbcow):

agreed...salman little help?

OpenStudy (anonymous):

its d^2f/dx dy the f(x)=xe^-8y+3y sorry chaps, this is the only info i have

OpenStudy (dumbcow):

ok, then assuming f(x) is not y, y is just another variable df/dx = e^-8y d/dy e^-8y = -8*e^-8y

OpenStudy (dumbcow):

for part a) df/dx = e^-7y d/dx e^-7y = 0